# ApisTox: a new benchmark dataset for the classification of small molecules toxicity on honey bees

### v1

@@ Abstract
The global decline in bee populations poses significant risks to agriculture, biodiversity, and environmental stability. To bridge the gap in existing data, we introduce ApisTox, a comprehensive dataset focusing on the toxicity of pesticides to honey bees (Apis mellifera). This dataset combines and leverages data from existing sources such as ECOTOX and PPDB, providing an extensive, consistent, and curated collection that surpasses the previous datasets. ApisTox incorporates a wide array of data, including toxicity levels for chemicals, details such as time of their publication in literature, and identifiers linking them to external chemical databases. This dataset may serve as an important tool for environmental and agricultural research, but also can support the development of policies and practices aimed at minimizing harm to bee populations. Finally, ApisTox offers a unique resource for benchmarking molecular property prediction methods on agrochemical compounds, facilitating advancements in both environmental science and cheminformatics. This makes it a valuable tool for both academic research and practical applications in bee conservation.
@@
Global declines of bee population are a serious threat to agricultural production, environmental stability, and overall biodiversity. Because bees are essential to the pollination of plants, protecting bee populations is of crucial importance for environmental preservation and food security. To evaluate the effects of different stressors, such as pesticides, infections, and environmental changes, on bee health, thorough and trustworthy data are needed. This necessity motivated the creation of ApisTox, a new state-of-the-art dataset on pesticide toxicity for honey bees (Apis mellifera), useful for the research and development community.
Existing datasets on bee toxicity, such as BeeTox [1] and subsets of ECOTOX [2], offer valuable insights but are limited by their scope, consistency, and the comprehensiveness of data. Furthermore, data on the effects of various crop management systems, as captured in CropCSM [3] datasets, are often disconnected from bee toxicity information, hindering holistic analyses. To address these gaps, we have developed a comprehensive dataset that combines carefully filtered and manually curated data from the above-mentioned resources. This integration makes it the largest and most consistent dataset on pesticide bee toxicity currently available.
The motivation behind creating this dataset stems from the need to understand the interactions between bees and their environments. The global decline in bee populations is a complex issue influenced by many factors, including exposure to pesticides and other human introduced variables, and climate change. By consolidating and curating data from multiple sources, our dataset allows for an assessment of a broader chemical space and provides a more coherent and comprehensive basis for analysis.
Our dataset encompasses a wide range of data types, including but not limited to, toxicity levels of various pesticides, herbicides, insecticides and other chemicals, with respect of their time of registration and references to external chemical databases and source publications. The data have been carefully filtered to ensure accuracy, relevance, and consistency, addressing a critical need for high-quality, standardized data in bee research.
@@
ApisTox is also one of the very few datasets outside medicinal chemistry that can be used for benchmarking molecular property prediction methods for graph classification. Predictive models utilizing molecular graphs are almost exclusively evaluated using data from de novo drug design, whereas agrochemical compounds possess quite different structural characteristics. Our dataset is also large enough for training and testing more data-demanding models, being larger than e.g. 17 datasets from Therapeutics Data Commons (TDC) benchmark [4]. ApisTox can serve as a part of challenging benchmarks for novel classification algorithms on molecular graphs, including out-of-distribution testing of models designed using medicinal chemistry data.
@@
The significance of this dataset extends beyond academic research. It can be useful for policy decisions and strategies, and for developing agricultural practices which minimize harm to bees. The dataset can help screen for bee-friendly chemicals and natural products, which help with the development of agricultural systems that promote bee health. Moreover, the dataset is designed with accessibility and interoperability in mind. This is intended to foster collaboration among researchers and encourage the development of novel, innovative solutions to protect bee populations. In conclusion, the creation of this new, state-of-the-art dataset for bee toxicity addresses a critical gap in existing data resources, offering a comprehensive, reliable, and accessible tool for studying the complex factors affecting bee health.
@@ Methods
We gather data about the acute honey bee toxicity of pesticides, based on their median lethal dose (LD50). We enrich it with additional metadata, which allows deeper analyses. Dataset is assembled from various sources, with extensive preprocessing, cleaning and deduplication. Those steps are described below in this section. The resulting files are described in Data Records section, with their characteristics and quality analyzed in Technical Validation section.
@@
ECOTOX [2] database is maintained by US EPA (United States Environmental Protection Agency) and consists of relatively raw data and experimental measurements. For each substance, it typically has many entries from different sources, with varied measurement values, and is the most comprehensive data source for ecotoxicology data. However, it provides relatively less structured data than some other databases, and in particular does not contain SMILES strings, typically processed by computational methods, instead relying on CAS (Chemical Abstracts Service) registry numbers. It is updated quarterly, and we use the version updated on 14th December 2023, which covers almost $1.2$ million measurements and over $13$ thousand chemicals.
Pesticide Properties DataBase (PPDB) [8] is a database that catalogs defined pesticide chemical entities, i.e. active ingredients, along with their physicochemical properties, ecotoxicological data, environmental fate, and human health impacts. It was created and is maintained by the Agriculture & Environment Research Unit (AERU) at the University of Hertfordshire. PPDB is a crucial source of curated, structured data for analyzing pesticide applications and facilitating risk management.
Bio-Pesticides DataBase (BPDB), also curated by the University of Hertfordshire, is our third source. It provides detailed information on the properties, efficacy, and application of pesticides originating from natural sources, such as microorganisms, plant extracts, and pheromones. BPDB entries typically comprise multiple components and lack a defined active ingredient.
Both the PPDB and BPDB are reviewed, managed, and updated through literature and legal resources, following established data curation protocols and guidelines. Those databases received endorsements from leading chemical and agrochemical organizations, including the International Union of Pure and Applied Chemistry (IUPAC) and the Food and Agriculture Organization (FAO). They are continuously updated, and we use data from 22nd February 2024.
All datasets undergo data processing and cleaning procedures, in order to unify their structure and fill missing fields. The entire workflow has been summarized in Figure 1. During the entire data cleaning and processing, we save all removed rows, along with removal reason, in a separate file, and include it along the rest of the raw and processed data (see Data Records section for details). Manual analysis of this data allowed us to verify the correctness of processing steps. In further sections, we describe them, and their rationale, in detail.
@@
ECOTOX database offers a rich set of filters, allowing flexible retrieval of data. We apply the following ones:
@@
    Endpoint “LD50” - endpoints are measurable numerical targets for given effects. We use median lethal dose (LD50), for consistency with PPDB and BPDB.
@@
We do not filter data by observed duration, because it contains measurements with at most 8 days, which still corresponds to acute toxicity, and most measurements are for 72 hours or less. The resulting set contains 2674 raw measurements, which then undergo a series of processing steps to obtain data aggregated per pesticide. For each measurement, we select substance name, CAS number, exposure type, observed response mean and unit. They identify the pesticide used, how it was applied, and measured LD50.
@@
Toxicity is typically divided into oral, contact, or other way of application, and PPDB and BPDB also use this system. However, ECOTOX applies a more fine-grained classification, and as a first step of processing, we map the specific exposure types into three standardized toxicity types as follows:
@@
  • “Multiple routes between application groups”, “Oral via capsule”, “Spray, unspecified”, “Environmental, unspecified” - other
To obtain a dataset with a single pesticide per row, raw measurements from ECOTOX need to be aggregated. EPA guidelines [9] suggest 11 $\mu\text{g / org}$ as the threshold for acute bee toxicity, i.e. LD50 at or below this value marks a pesticide definitely toxic for honey bees. PPDB and BPDB [10], for consistency with EU and UK regulatory institutions, use 1 $\mu\text{g / org}$ and 100 $\mu\text{g / org}$ as thresholds, i.e. LD50 equal to or lower than 1 means highly toxic pesticide, between 1 and 100 moderately toxic, and higher than 100 means non-toxic or slightly toxic substance. We therefore calculate two such labels for each pesticide: binary non-toxic/toxic (EPA label), and ternary non-toxic/moderately toxic/highly toxic (PPDB level). They are numerically encoded as 0/1 and 0/1/2, respectively.
We identify each distinct substance by the CAS number, and for each one we calculate two labels for every toxicity type (oral, contact, other) available for that pesticide. Those measurements vary strongly for some pesticides, so we take a conservative approach here, to ensure high data quality. If the lowest and highest measurement agree for the EPA binary label, i.e. all measurements are below or above 11 $\mu\text{g / org}$, we assign positive or negative class, respectively. Otherwise, we mark the toxicity label as “Unspecified”, and remove such rows. Then we assign PPDB ternary level based on median measurement. Lastly, for each pesticide, we take the strongest toxicity type (e.g. with the highest level), as representing the most toxic effect the pesticide can have on bees. We also save the information which toxicity type was the strongest for each pesticide.
@@
Lastly, we add information about agrochemical type for each substance. In PubChem, compounds can have a separate section “Agrochemical information” with descriptions of agricultural applications. Using PUG REST API, we fetch descriptions from this section for each substance, and we search for keywords “herbicide”, “fungicide” and “insecticide”. If any of those words appear in the description, we note that information as agrochemical type. If none are found, but the compound has this section, it means that it has other agrochemical applications, e.g. as a growth agent or fertilizer, and we mark it as “other agrochemical”. This creates a total of four boolean (binary) variables. In the last case, when the compound page does not have “Agrochemical information” section at all, it can have all zeros, meaning “unknown” pesticide type. We still include those substances, since we have bee toxicity measurements for them. They could have been researched as potential pesticides, but not yet registered or used.
@@
Both PPDB and BPDB databases have the same structure organized in tables, available via web pages. We extract all relevant values using regular expressions on HTML responses. This approach always utilizes the latest available data. We downloaded the data on 22nd February 2024.
For each pesticide, LD50 values for oral, contact, and other honey bee toxicity can be provided. Often no measurement, or only one or two are available, and we ignore substances without any toxicity measurement for honey bees. If we have more than one measurement, we take the lowest value, i.e. the strongest toxicity. Additionally, some values are provided in a non-standard format, especially for low toxicity, and marked as “Low”, “Non-toxic”, or in scientific notation, e.g. “10^3”. We normalize those cases as non-toxic, i.e. label and level 0.
@@
CAS numbers and SMILES strings are available for almost all pesticides, and we exclude ones without this information. CID numbers are also often available, but in case they are missing, we fill them using PubChem and PUG REST API, based on CAS and SMILES.
Combining datasets
@@
After gathering three preprocessed datasets, we merge them. Firstly, we concatenate all rows and drop obvious duplicates, i.e. compounds with the same CAS, SMILES and label.
@@
We ensure that the dataset consists of only distinct molecules. All SMILES are canonicalized using RDKit 2023.9.5 [13], and this way we also make sure that all SMILES strings are valid and processable by this software. We then drop all duplicates, first using SMILES, and then using CAS. When removing duplicates, we keep rows from datasets in the order of preference: PPDB, BPDB, ECOTOX. This was motivated by the fact that PPDB and BPDB are additionally manually verified, and therefore can have slightly higher quality.
@@
Lastly, we add the first publication year information to all molecules, using their CID numbers and PubChem literature records. While this is not exact information about first usage of a given substance as a pesticide, it still provides a good approximation of when a given compound was first created. We noticed that PubChem has three obvious errors in this regard, attributing an unreasonably large number of molecules to three publications [14, 15, 16]. We verified manually that this is a mistake, and we exclude those papers from this mapping. We sort the resulting dataset by this year. It contains a total number of 1035 molecules.
@@
Statistical models, especially predictive machine learning (ML) models, require separate parts of data for fitting (training) models, and for their testing and verification of performance. For molecular graphs and chemical data especially, there are many non-obvious possible sources of data leakage, which lead to improper and overly optimistic estimation of models’ performance. Having predetermined splits is beneficial for reproducible science in ML, therefore we compute train-test splits and distribute those files along the full dataset.
@@
We do not provide scaffold split [17], popular in medicinal chemistry, for the following reasons. It uses the Bemis-Murcko scaffolds to group molecules, and then puts the smallest groups in the test set. The idea is to put the most structurally different compounds into the test set, which requires generalization to new ares of chemical space. However, this assumes that those scaffolds can be calculated at all, and that the dataset will contain many small groups of scaffolds. Bemis-Murcko scaffolds are not defined for ring-free compounds, as well as those with disconnected components (e.g. salts), which constitute almost 20% of the data (see Technical Validation section for more details). This results in one large “no scaffold” group in the training set, therefore this approach does not differentiate them structurally. In the worst case, almost identical molecules can be both in training and test set, introducing data leakage. This is clearly a problem, and for this reason we recommend using MaxMin or time split instead.
Stratified random split puts data points randomly into training and testing parts, disregarding the features or internal structure of molecular graphs. Stratification ensures that the proportion of toxic and non-toxic classes is approximately the same in the full dataset and both splits. This is desirable, since ApisTox is imbalanced and toxic molecules constitute a minority class. We use binary toxicity labels here. This split is susceptible to clustering in chemical space, and therefore can result in structurally similar pesticides in both training and testing sets [18, 17]. Depending on application, this may be seen as a form of data leakage and artificially increase test score.
Time split puts the newest molecules in the testing set, in order to simulate the actual discovery and adoption of pesticides. The underlying assumption is that intrinsically new substances are introduced over time. This is often a very realistic setting, especially for designing novel molecules, but this information is rarely available in molecular datasets [18, 19]. Our literature-based year assignment, while not perfectly precise, allows using this kind of split.
@@
We use Scikit-learn [22] for computing stratified random split, Pandas [23] for time split, and DeepChem [24] for MaxMin split (with Dice distance and 1024 bits ECFP4 fingerprints).

### v3

@@ Abstract
The global decline in bee populations poses significant risks to agriculture, biodiversity, and environmental stability. To bridge the gap in existing data, we introduce ApisTox, a comprehensive dataset focusing on the toxicity of pesticides to honey bees (Apis mellifera). This dataset combines and leverages data from existing sources such as ECOTOX and PPDB, providing an extensive, consistent, and curated collection that surpasses the previous datasets. ApisTox incorporates a wide array of data, including toxicity levels for chemicals, details such as time of their publication in literature, and identifiers linking them to external chemical databases. This dataset may serve as an important tool for environmental and agricultural research, but also can support the development of policies and practices aimed at minimizing harm to bee populations. Finally, ApisTox offers a unique resource for benchmarking molecular property prediction methods on agrochemical compounds, facilitating advancements in both environmental science and chemoinformatics. This makes it a valuable tool for both academic research and practical applications in bee conservation.
@@
Bees are essential to the pollination of plants, and protecting their populations is of crucial importance for environmental preservation and food security [1, 2]. Global declines of their population are a serious threat to agricultural production, environmental stability, and overall biodiversity [3]. It is a complex issue influenced by many factors, including exposure to pesticides, other human introduced variables, and climate change [4]. However, data on the effects of various crop management systems, as captured e.g. in CropCSM [5] datasets, are often disconnected from bee toxicity information, hindering more holistic analyses. To evaluate the effects of different stressors, such as pesticides, infections, and environmental changes on bee health, thorough and trustworthy data is needed [6, 7]. This necessity motivated the creation of ApisTox, the largest single resource on honey bee (Apis mellifera) toxicity data publicly available.
None of the currently available databases cover the full range of pesticides and other toxic substances that bees are exposed to in real-world scenarios. In fact, bee toxicity data is scattered around many sources, encompassing both specialized ones (e.g. BeeTox [8]), and general-purpose pesticide or ecotoxicology databases, such as PPDB [9], BPDB [9], and ECOTOX [10]. Existing datasets on bee toxicity offer valuable insights, but are hindered by many inconsistency problems. For example, toxicity data points exist in different formats, units, and organization across reference databases. Databases also frequently list substances not relevant to honey bees, requiring data filtering and curation. Apart from the data quality, problems occur also with records duplication or improper chemical information, which can even be unreadable for chemoinformatics programs and pipelines. Lack of additional data relevant for QSAR, such as formulation use or pesticide introduction date, is further limiting the scope, consistency, and comprehensiveness of currently available bee toxicity data.
To address some of these gaps, we have developed a comprehensive dataset that combines carefully filtered and manually curated data from the above-mentioned resources and augment it with additional relevant information. By consolidating and curating data from multiple sources, our dataset allows for an assessment of a broader chemical space and provides a more consistent basis for further analysis. ApisTox encompasses a range of information types, including toxicity levels of four classes of agents: herbicides, fungicides, insecticides and other agrochemicals. All chemical records are additionally augmented with their date of first occurrence in the literature and references to the source chemical database. The data has been rigorously filtered and curated, to ensure accuracy and consistency, addressing a critical need for high-quality, standardized data in bee research.
@@
ApisTox is also one of the very few datasets outside medicinal chemistry that can be used for benchmarking molecular property prediction methods. Currently, machine learning (ML) models are almost exclusively evaluated using data derived from the medicinal fields, whereas agrochemical compounds possess quite different structural and physicochemical characteristics. Our dataset is large enough for training and testing more data-demanding models, being larger than e.g. 17 datasets from Therapeutics Data Commons (TDC) benchmark [11]. We also provide explicit train-test splits, including maximum diversity (MaxMin) split and a novel approximation of time split using PubChem literature data. This ensures realistic, challenging evaluation and fair comparison of algorithms. Therefore, ApisTox can serve as a part of benchmarks for novel classification algorithms, including out-of-distribution testing of models designed using medicinal chemistry data.
@@
Finally, the significance of this dataset extends beyond academic research. The dataset can help screen for bee-friendly chemicals and natural products, which help with the development of agricultural systems that promote bee health [12, 13]. It can also be useful for policy decisions and strategies, particularly for evidence based decisions, e.g. assessing the impact of changing regulations and R&D funding on the number of first-in-class type of compounds developed over time. ApisTox is designed with accessibility and interoperability in mind. This is intended to foster collaboration among researchers and encourage the development of novel, innovative solutions to protect bee populations. In conclusion, the creation of this new, high quality dataset for bee toxicity addresses a critical gap in existing data resources, offering a comprehensive, consistent, and accessible tool for studying the complex factors affecting bee health.
@@ Methods
We gather data about the acute honey bee toxicity of pesticides. Following the US EPA guidelines [14], as well as EU and UK regulatory standards [15, 16], we focus on the median lethal dose ($LD_{50}$) as a toxicity measure. The same guidelines designate honey bees as a major organism for regulatory decisions, due to their ecological and economic importance.
We enrich it with additional metadata, which allows deeper analyses. Dataset is assembled from various sources, with extensive preprocessing, cleaning and deduplication. The resulting files are described in the “Data Records” section, with their characteristics and quality analyzed in the “Technical Validation” section.
@@
ECOTOX [10] (https://cfpub.epa.gov/ecotox/) database is maintained by US EPA (United States Environmental Protection Agency) and consists of relatively raw data and experimental measurements. For each substance, it typically has many entries from different sources, with varied measurement values, and is the most comprehensive data source for ecotoxicology data. However, it contains relatively less structured data than some other databases, and in particular does not provide SMILES strings, typically processed by computational methods. Instead, it relies on CAS (Chemical Abstracts Service) registry numbers. It is updated quarterly, and we use the version updated on 14th December 2023, which covers almost $1.2$ million measurements and over $13$ thousand chemicals.
Pesticide Properties DataBase (PPDB) [9] (https://sitem.herts.ac.uk/aeru/ppdb/en/index.htm) is a database that catalogs defined pesticide chemical entities, i.e. active ingredients, along with their physicochemical properties, ecotoxicological data, environmental fate, and human health impacts. It was created and is maintained by the Agriculture & Environment Research Unit (AERU) at the University of Hertfordshire. PPDB is a crucial source of curated, structured data for analyzing pesticide applications and facilitating risk management.
Bio-Pesticides DataBase (BPDB) (https://sitem.herts.ac.uk/aeru/bpdb/index.htm), also curated by the University of Hertfordshire, is our third source. It provides detailed information on the properties, efficacy, and application of pesticides originating from natural sources, such as microorganisms, plant extracts, and pheromones. BPDB entries often comprise multiple components and lack a defined active ingredient.
Both the PPDB and BPDB are reviewed, managed, and updated through literature and legal resources, following established data curation protocols and guidelines. Those databases received endorsements from leading chemical and agrochemical organizations, including the International Union of Pure and Applied Chemistry (IUPAC) and the Food and Agriculture Organization (FAO). They are continuously updated, and we use data obtained at 22nd February 2024. Raw data obtained at this date is included with other files for reproducibility, as described in detail in section “PPDB and BPDB”.
All datasets undergo data processing and cleaning procedures, in order to unify their structure and fill missing fields. The entire workflow has been summarized in Figure 1. The workflow itself constitutes a major novel contribution in this work, fully describing the set of procedures for parsing, merging, cleaning and deduplication of those three databases, with the reusable, open source code available.
During the entire data cleaning and processing, we save all removed rows, along with removal reason, in a separate file, and include it along the rest of the raw and processed data (see “Data Records” section for details). Manual analysis of this data allowed us to verify the correctness of processing steps. In further sections, we describe them, and their rationale, in detail.
@@
Classification labels
For all three databases, we apply the same procedure for creating the classification labels of honey bee toxicity, based on $LD_{50}$ values. As such, ApisTox can be used for classification, rather than regression, and here we explain the rationale of this decision.
Toxicity studies, in particular QSAR models, often use classification approach, rather than regression. Examples include Tox21 and ToxCast datasets from MoleculeNet [20], as well as most toxicity datasets from TDC [11]. If direct measurement values are available, most often $LD_{50}$ or $LC_{50}$, they can be binarized into toxic/non-toxic classes, using some predefined threshold. Sometimes more classes are used, e.g. ternary toxic / moderately toxic / highly toxic. Such approach has also been used e.g. for rats [21, 22], fish and Daphnia Magna [23], and earthworms [24].
Such approach is used particularly due to the inherent uncertainty of toxicity measurements, in particular for $LD_{50}$. It is a statistically obtained virtual value, constituting a best-effort estimation of the dose required to kill 50% of animals [25]. There are many different protocols for calculating $LD_{50}$, differing in complexity and quality [25, 26]. For experiments on rats, $LD_{50}$ values can vary as much as 3- to 11-fold between different laboratories [27], and similarly significant differences have been observed for honey bees [28]. For low toxicity pesticides (with $LD_{50}$ values greater or equal to $100$ $\mu\text{g / bee}$), measuring $LD_{50}$ for honey bees is particularly problematic, and EFSA guidelines [16] contain dedicated procedures and limit tests for those cases. As such, the mean or median aggregate value for regression would itself be highly uncertain. This is also why PPDB and BPDB often don’t provide exact $LD_{50}$ values, but rather a range limit, e.g. “>300”. In classification, we only require the estimated $LD_{50}$ values to lie below or above a given threshold, lessening the impact of measurement errors.
Furthermore, in ecotoxicology regulatory organs recognize this issue and provide the official $LD_{50}$ thresholds for different organisms. In practice, when designing new pesticides, one is not necessarily interested in the exact $LD_{50}$ value but rather if the new substance will be accepted by the regulatory organ, i.e. if the toxicity is low enough. Classification models explicitly provide this information, when an appropriate threshold is used, with predicted probability giving a measure of certainty in the prediction.
US EPA guidelines [14] for honey bees use binary toxic/non-toxic labels, with $11$ $\mu\text{g / org}$ as the threshold for acute bee toxicity. This means that $LD_{50}$ less or equal to this value is deemed toxic to honey bees. We mark toxic class as $1$, and non-toxic as 00, as shown in Equation 1.
$label(LD_{50})=\begin{cases}1&\text{if}\quad LD_{50}\leq 11\\ 0&\text{if}\quad LD_{50}>11\end{cases}$ (1)
PPDB, based on EU and UK regulatory guidelines [15, 16], introduced a 3-level (ternary) highly toxic/moderately toxic/non-toxic labels, with thresholds $1$ $\mu\text{g / org}$ and $100$ $\mu\text{g / org}$. We mark highly toxic as $2$, moderately toxic as $1$, and non-toxic as 00, as shown in Equation 2.
$ppdb\_level(LD_{50})=\begin{cases}2&\text{if}\quad LD_{50}\leq 1\\ 1&\text{if}\quad 1<LD_{50}\leq 100\\ 0&\text{if}\quad LD_{50}>100\end{cases}$ (2)
To maximize the applicability of ApisTox, we calculate classes using both of those approaches: binary label and ternary ppdb_level. As such, it can be used either for binary classification or ordinal classification with three levels.
@@
ECOTOX database offers a rich set of filters, allowing flexible retrieval of data. It is divided into “Aquatic” and “Terrestial” organisms. We select the latter, and apply the following filters:
@@
    Endpoint “LD50” - endpoints are measurable numerical targets for given effects. We use median lethal dose ($LD_{50}$), for consistency with PPDB and BPDB, following EPA guidelines [14] and EU and UK regulatory standards [15, 16].
@@
We do not filter data by observed duration, because it contains measurements with at most 8 days, which still corresponds to acute toxicity, and most measurements are for 72 hours or less. The resulting set contains 2674 raw measurements, which then undergo a series of processing steps to obtain data aggregated per pesticide. For each measurement, we select substance name, CAS number, exposure type, observed response mean and unit. They identify the pesticide used, how it was applied, and measured $LD_{50}$.
@@
Toxicity is typically divided into oral, contact, or other way of application. PPDB and BPDB also use this system. However, ECOTOX applies a more fine-grained classification, so we map the ECOTOX exposure types into three standardized toxicity types as follows:
@@
  • “Multiple routes between application groups”, “Oral via capsule” [^1] Only a single row had the “Oral via capsule” value, and this was not a standard oral diet in the original experimental study[29], so we include it as “other”, rather than “oral”., “Spray, unspecified”, “Environmental, unspecified” - other
To obtain a dataset with a single pesticide per row, raw $LD_{50}$ measurements from ECOTOX need to be aggregated into a single number. We identify each distinct substance by the CAS number. They can have $LD_{50}$ measurements for oral, contact, and other toxicity type. For many pesticides, only one or two of those are available. Their values vary strongly for some pesticides, so we take a conservative approach here, to ensure high data quality.
For each toxicity type, we check if the lowest and highest measurement agree for the EPA guidelines threshold, i.e. if all values are below or above $11$ $\mu\text{g / org}$. If so, we assign toxic or non-toxic class, respectively. Otherwise, we mark the toxicity label as “Unspecified”, and remove such rows. This way, we further alleviate problems arising from $LD_{50}$ measurement uncertainty, as described in the “Classification labels” section.
We also assign PPDB ternary level based on median measurement value. Lastly, for each pesticide, we take the strongest toxicity type, as representing the most toxic effect the pesticide can have on bees. For example, if it has oral toxicity $2$ (highly toxic) and contact toxicity $1$ (moderately toxic), we mark it as $2$. This way, only pesticides truly safe for honey bees will get level 00. We also save the information which toxicity type was the strongest for each pesticide.
@@
Lastly, we add information about agrochemical type for each substance, i.e. whether it is a herbicide, fungicide, insecticide, or other agrochemical. This allows users to carry out more detailed analyses, e.g. narrow down toxicity studies only to insectides, which are typically the most dangerous pesticide type for bees. In addition, predictive models and their performance can be analyzed for those types. For example, one can check if the model performs equally well for fungicides and insecticides.
In PubChem, compounds can have a separate section “Agrochemical information” with descriptions of agricultural applications. Using PUG REST API, we fetch descriptions from this section for each substance, and we search for keywords “herbicide”, “fungicide” and “insecticide”. If any of those words appear in the description, we note that information as agrochemical type. If none are found, but the compound has this section, it means that it has other agrochemical applications, e.g. as a growth agent or fertilizer, and we mark it as “other agrochemical”. This creates a total of four boolean (binary) variables. If the compound page does not have “Agrochemical information” section, all four type features are zeros, meaning “unknown” pesticide type. We still include those substances, since we have bee toxicity measurements for them. They could have been researched as potential pesticides, but not yet registered or widely used in agriculture.
@@
PPDB and BPDB databases have similar structure, organized in HTML tables, available via web pages. They contain a very comprehensive set of information about pesticides (over 300 different fields). We extract only some fields, from which we compute relevant features using regular expressions on HTML responses, as described below. All molecules that do not have toxicity measurements for honey bees are ignored. This approach always utilizes the latest available data, and we use the data obtained on 22nd February 2024.
For each pesticide, $LD_{50}$ values for oral, contact, and other honey bee toxicity can be provided. If we have more than one toxicity type, we take the lowest value, i.e. the strongest toxicity. We create binary label and ternary level, as described in the “Classification labels” section, similarly to ECOTOX processing. Additionally, some values are provided in a non-standard format, especially for low toxicity, and marked as “Low”, “Non-toxic”, or in scientific notation, e.g. “10^3”. We normalize those cases as non-toxic, i.e. label and level 0.
@@
CAS numbers and SMILES strings are available for almost all pesticides, and we exclude those without this information. CID numbers are also often available, but in case they are missing, we fill them using PubChem and PUG REST API, based on CAS and SMILES.
Molecular standardization and deduplication
@@
After gathering three preprocessed datasets, we merge them. Firstly, we concatenate all rows and drop obvious duplicates, i.e. compounds with the same CAS, SMILES and label. However, since SMILES strings are not unequivocal, this does not yet fully ensure lack of structural duplicates among molecules. Thus, we standardize molecules with RDKit and create canonical SMILES, which can be used for more reliable deduplication.
We perform multi-step canonicalization and sanitization based on input SMILES, using RDKit 2023.9.5 [32]. This way, we make sure that all SMILES strings are valid and processable by this software, and all molecules conform to basic sanity checks like proper valence, all bonds are represented in a unified way etc. This is done with MolFromSmiles function, which calls a whole sanitization pipeline underneath [33], and with MolToSmiles function, which generates canonical SMILES. This pipeline includes:
  • valence cleanup, handling non-standard valence states
  • kekulization
  • determining aromaticity of bonds and atoms
  • calculation of bond conjugation and atom hybridization states
  • chirality cleanup, removing chiral tags from atoms that are not sp3 hybridized
  • adjusting hydrogens, leaving only necessary explicit ones (e.g. for heteroatoms in aromatic rings)
Those steps are common for most molecular standardization workflows [34, 35, 36]. We perform quite conservative set of operations, to allow end users to use their own more aggressive transformations or filtering if desired. In particular, following EPA guidelines for pesticide QSAR (created in conjunction with NAFTA Technical Working Group on Pesticides) [37], we do not remove salts, inorganics, mixtures, or metal-containing compounds. This is particularly relevant for toxicity studies of organic salts [38]. Most standardization workflows focus solely on medicinal chemistry, where such molecules are commonly removed, but this would get rid of much important information prevalent in agrochemistry. For example, Mancozeb and Maneb are treated as independent pesticides in regulatory issues [39], while they only differ in the complexation with metal ions (Mancozeb is complexed with manganese and zinc ions, Maneb is complexed only with manganese ions).
@@
Such standardization ensures that all agrochemicals in ApisTox are processable by RDKit, which is the most commonly used open source chemoinformatics tool in Python, and also integrates with e.g. KNIME. Resulting SMILES strings are canonical, and can be used to perform further deduplication. We then drop all duplicates, first using SMILES, and then using CAS. When removing duplicates, we keep rows from datasets in the order of preference: PPDB, BPDB, ECOTOX. This was motivated by the fact that PPDB and BPDB are additionally manually verified, and therefore can have slightly higher quality.
Adding PubChem literature date
Lastly, we add the first publication year information to all molecules, using their CID numbers and PubChem literature records. While this is not exact information about first usage of a given substance as a pesticide, it provides a good approximation of when a given compound has entered the public domain. With this information, additional analyses can be carried out with ApisTox, related to the development of pesticide structures through time and broader analysis of the effectiveness of regulations and R&D spendings. It can also be directly used in molecular property prediction to create a time split (also known as chronological split) for train-test sets. This approximation should be much more precise than scaffold split [20] or SIMPD [40]. Those techniques were created to approximate time split, but are based on some predefined structural biases, e.g. Bemis-Murcko scaffolds, which PubChem dates do not enforce.
@@
We noticed that PubChem has three obvious errors in this regard, attributing an unreasonably large number of molecules to three publications [41, 42, 43]. We verified manually that this is a mistake, and we exclude those papers from this mapping. We sort the resulting dataset by this year. It contains a total number of 1035 molecules.
@@
Statistical models, especially predictive machine learning (ML) models, require separate parts of data for fitting (training) models, and for their testing and verification of performance. For molecular graphs and chemical data especially, there are many non-obvious possible sources of data leakage, which lead to improper and overly optimistic estimation of models’ performance. Having predetermined splits is beneficial for reproducible science in ML, therefore we compute train-test splits and distribute those files along the full dataset. To the best of our knowledge, ApisTox is the first dataset of honey bee toxicity to provide explicit train-test splits and evaluation protocol, making it better suited for comparison of ML techniques.
@@
We do not provide scaffold split [20], popular in medicinal chemistry, for the following reasons. It uses the Bemis-Murcko scaffolds to group molecules, and then puts the smallest groups in the test set. The idea is to put the most structurally different compounds into the test set, which requires generalization to new areas of chemical space. This way, it aims to approximate a time split, with an assumption that future compounds have new scaffolds. However, this assumes that those scaffolds are structurally meaningful, and that the dataset will contain many small groups of scaffolds. Bemis-Murcko scaffolds are not defined for ring-free compounds, or for those with disconnected components (e.g. salts), since their definition relies on connected ring systems [44]. Such molecules constitute almost 20% of the data (see the “Technical Validation” section for more details), and that group would result in one large “no scaffold” group in the training set. Therefore, this approach does not differentiate them structurally and, in the worst case, almost identical molecules can be in training and test sets, introducing data leakage. This is clearly a problem, and for this reason we recommend using MaxMin or time split instead. See also [45] for other potential caveats with scaffold-based approaches.
Stratified random split puts data points randomly into training and testing parts, disregarding the features or internal structure of molecular graphs. Stratification ensures that the proportion of toxic and non-toxic classes is approximately the same in the full dataset and both splits. This is desirable, since ApisTox is imbalanced and toxic molecules constitute a minority class. We use binary toxicity labels here. Such split is susceptible to clustering in chemical space, and therefore can result in structurally similar pesticides in both training and testing sets [46, 20]. Depending on application, this may be seen as a form of data leakage and artificially increase test score.
Time split puts the newest molecules in the testing set, in order to simulate the actual discovery and adoption of pesticides. The underlying assumption is that intrinsically new substances are introduced over time. This is often a very realistic setting, especially for designing novel molecules, but such information is rarely available in molecular datasets [46, 47]. Our literature-based year assignment, while not perfectly precise, allows using this kind of split. Using PubChem to approximate the time split is, as far as we are aware, a novel approach, and a meaningful alternative to e.g. scaffold split or SIMPD method [40].
@@
We use Scikit-learn [50] for computing stratified random split, Pandas [51] for time split, and DeepChem [52] for MaxMin split (with default Dice distance and 1024 bits ECFP4 fingerprints).